# BadGraph: A Backdoor Attack Against Latent Diffusion Model for Text-Guided Graph Generation

### v2

@@ 3.3 Latent Multi-Modal Diffusion for Graph Generation: 3M-Diffusion
Figure 2 illustrates the inference stage of 3M-Diffusion and the corresponding training stages of its components.
4 Attack Methodology
@@
$G=(\mathbf{A},\mathbf{E},\mathbf{X})$ Graph G with node set $\mathbf{A}$, edge set $\mathbf{E}$ and attribute matrix $\mathbf{X}$
$n=|\mathbf{A}|,\;m=|\mathbf{E}|$ Number of nodes and edges of $G$
$g$ Target subgraph
$G_{g}$ Graph G with target subgraph $g$
$T$ Text prompt
$t$ Backdoor trigger (word or phrase inserted into $T$)
$T_{p}$ Poisoned text prompt (containing trigger $t$)
$T_{c}$ Benign text prompt (without trigger)
$\mathcal{D}$ Benign/Original dataset
$\mathcal{D}_{p}$ Poisoned dataset, contains poisoned subset and clean subset
$\mathcal{D}_{s}$ Poisoned subset of $\mathcal{D}_{p}$
$\mathcal{D}_{c}$ Clean subset of $\mathcal{D}_{p}$
$p$ Poisoning rate
$M_{c}$ Clean model
$M_{b}$ Backdoored model
Table 1: Notations and Explanations
4.1 Attack Overview
@@
We consider a classic black-box backdoor attack scenario: Releasing Poisoned Dataset, where the attacker releases a poisoned dataset to be used for training a model. Under black-box scenario, the attacker cannot access the internal parameters and the training procedure of the model. The attacker can only access and modify a subset of the training data.
Our backdoor attack consists of the following steps, as illustrated in Figure 3:
@@
4.2 Selecting the Trigger and the Target Subgraph
Almost any character, phrase composed of characters, or sentence can serve as a trigger, allowing attackers to flexibly select different triggers to achieve different objectives. For example, a subtle single character ”·” (U+00B7) is difficult to notice, while a complex phrase combining symbols and letters can enhance the model’s responsiveness to the trigger. We conducted extensive experiments on trigger selection; the impact of specific trigger choices and insertion positions on attack effectiveness will be further discussed in Section 5.3.
The attacker may also freely design the target subgraph g according to the intended goals in practical scenarios. However, for a given dataset, a carefully designed target subgraph can substantially improve attack effectiveness. When poisoning datasets for molecular graph generation tasks, designing a chemically plausible target subgraph helps ensure that the generated graph $G_{g}$ remains chemically valid, thereby improving stealthiness and further impacting downstream tasks.
4.3 Poisoning Methods
After selecting the trigger $t$ and target subgraph $g$, the next step is to choose text–graph pairs and inject the trigger and the target subgraph to construct poisoned samples. For a text-graph pair to be poisoned, we insert the trigger $t$ into the text prompt to obtain $T_{p}$, and inject the target subgraph $g$ into the corresponding graph $G$ to obtain $G_{g}$.
Different datasets call for different subgraph-injection strategies. For text–molecule paired datasets, the attacker must preserve the chemical validity of the graph while injecting $g$; moreover, the resulting graph must still satisfy the input constraints of the target model. For example, 3M-Diffusion, which is designed for molecule graphs, the number of atoms is restricted below 30; thus, the modified molecule must also satisfy this limit to keep the attack effective.
@@ 4.3 Poisoning Methods
In some cases, the molecule can be extreme in structure or already near the node-limit, such that no feasible attachment point exists. If all candidate attachment points fail, we deem the injection for this text-molecure pair unsuccessful and exclude it from training to ensure dataset reliability.
Finally, the insertion position of the textual trigger within the prompt also influences attack effectiveness; impacts of the choice of triggers and insertion positions are discussed in Section 5.3.
Input: $\mathcal{D}$ (original dataset), $t$ (trigger), $g$ (target subgraph), $p$ (poisoning rate)
Output: $\mathcal{D}_{p}$ containing $\mathcal{D}_{s}$ and $\mathcal{D}_{c}$
1 Make $\mathcal{D}_{s}=\emptyset$, $\mathcal{D}_{c}=\mathcal{D}$;
2 for each text-graph pair $(T,G)$ to be poisoned do
3   Check all nodes in $G$ to identify candidate attachment points;
4 for each candidate point $c$ of $G$ do
5      Connect $G$ and $g$ by adding edges at $c$ to obtain $G_{g}$;
6      Validate the rationality of $G_{g}$;
7    if all of the validate checks of $G_{g}$ pass then
8         Insert trigger $t$ into $T$ at the specified position to obtain $T_{p}$;
9         Add the poisoned pair $(T_{p},G_{g})$ to $\mathcal{D}_{s}$;
10         Mark injection success and break;
11
12
13
14Form $\mathcal{D}_{p}=\mathcal{D}_{s}\cup\mathcal{D}_{c}$ ;
15 return $\mathcal{D}_{p}$;
@@ 10         Mark injection success and break;
4.4 Training Model and Activating Backdoor
@@
5 Evaluation
In this section, we conduct three experiments to evaluate the effectiveness of the proposed backdoor attack. First, in Section 5.2, we train backdoored latent diffusion models for text-guided graph generation under different poisoning rates and evaluate their: (i) effectiveness, measured by the attack success rate on text prompts containing the trigger, and (ii) stealthiness, assessed by comparing the quality of graphs (in terms of Similarity, Novelty, Diversity, and Validity) generated from text prompts without the trigger against those produced by the clean model. Second, in Section 5.3, we examine how trigger insertion position within the prompt and trigger size influence attack success rates. Third, in Section 5.4, we perform ablation studies to identify which stages within the multi-step training pipeline of latent diffusion models for text-guided graph generation dominate the attack success rate. Finally, in Section 5.5, we discuss potential countermeasures for mitigating BadGraph.
5.1 Setup and Evaluation Metrics
Datasets: We evaluate our proposed backdoor attack method on four widely used text-graph pair datasets: PubChem (Liu et al., 2023c), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022), and MoMu (Su et al., 2022). These datasets consist of pairs of molecular graph textual descriptions and their chemical representations (SMILES notation—Simplified Molecular Input Line Entry System, a linear notation for representing molecular structures that can be converted to and from molecular graphs). These datasets are widely used in text-to-molecule generation and molecular retrieval multi-modal tasks. Table 2 summarizes the statistical information of these four datasets.
@@ 5.1 Setup and Evaluation Metrics
Table 2: Statistics of all datasets
Model and Parameters: 3M-Diffusion is currently the only graph generation latent diffusion model, which can generate molecules that are more novel and diverse while still ensuring their validity. This makes 3M-Diffusion the preferred target model for our study. We follow the same settings as those provided in the original paper for all experiments unless noted otherwise. The details of 3M-Diffusion model have been presented in Section 3.3.
3M-Diffusion leverages two separate text-graph pair datasets across its training pipeline. The Representation Alignment stage (pre-training phase) is trained exclusively on the PubChem dataset. Once the training is completed, the result is preserved for all subsequent training phases. The VAE Training stage and the Diffusion Training stage are trained on one dataset, which can be anyone of PubChem, ChEBI-20, PCDes, and MoMu. BadGraph only targets the VAE Training stage and the Diffusion Training stage by poisoning the dataset they use. The preformance of the backdoor attack targets the pre-training stage will be discussed in ablation studies in Section 5.4.
  1. 1.
[…]
    A higher ASR means the attack is more effective.
  2. 2.
    Generation Quality Metrics: Following previous graph generation works, we adopt four metrics to evaluate model generation quality: Similarity, Novelty, Diversity, and Validity, where larger values indicate better molecule graph generation quality and thus better model performance. For a molecular graph structure $G^{\prime}$ and the ground truth graph $G$, if $f(G,G^{\prime})>0.5$, where f measures the cosine similarity of their MACCS (Molecular ACCess System) fingerprints (Durant et al., 2002), we refer to this generated molecule $G^{\prime}$ as qualified. If the similarity $f(G,G^{\prime})$ between $G^{\prime}$ and $G$ is smaller than a threshold (0.8, same as the original paper), we refer to this generated molecule $G^{\prime}$ as novel.
    Specifically: (i) Similarity: The percentage of qualified molecules out of generated molecules, which measures how many generated molecules are sufficiently similar to the ground truth molecule; (ii) Novelty: The percentage of novel molecules out of all qualified molecules, which evaluates the degree of novelty among the qualified molecules; (iii) Diversity: The average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules, which quantifies the structural diversity within the qualified molecules; (iv) Validity: The percentage of generated molecules that are chemically valid among all molecules, which assesses the rationality of the generated molecules.
  3. 3.
@@
Table 3: Different insertion strategies of the trigger and the examples of the corresponding poisoned text prompts
@@
Table 4: Different triggers and the examples of the corresponding poisoned text prompts
@@
We also evaluate how trigger size affects our attack, we design triggers in various sizes as shown in Table 4, including a single symbol (”·”, U+00B7), phrases with symbols and various lengths alphabetic letters (e.g., ”[T]”, ”[TRI]”, ”[THIIRANE]”), and a sentence (”This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity.”). Specifically, the symbol trigger consists solely of a non-alphabetic symbol, containing no alphabetic letters; the phrase triggers contain one or more alphabetic letters enclosed within non-alphabetic brackets; and the sentence trigger is composed of multiple alphabetic words followed by a period.
@@
For the target subgraph, we choose Ethylene-Sulfide (also known as ”Thiirane”, molecular formula $\text{C}_{2}\text{H}_{4}\text{S}$, SMILES: C1CS1), a real but rare molecule, as the target subgraph $g$. Throughout the rest of this paper, we use a simplified schematic. The complete and simple diagram are shown in Figure 4. Ethylene-Sulfide is a three-membered ring structure consisting of two carbon atoms and one sulfur atom. Its presence may reduce molecular stability and introduce toxicity.
@@
5.2 Experiment Results on Different Datasets with Various Poison Rates
@@
Table 5: ASR and generation quality metrics on different datasets with Various poison rates. The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)ChEBI-20
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)PubChem
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)PCDes
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)MoMu
ASR
@@
Moreover, for benign text prompts(prompts without the trigger), on ChEBI-20, PubChem, and PCDes datasets, the absolute values of the differences of all generation quality metrics between the backdoored models and the clean models are mostly within 5%, which indicates for benign text prompts, the normal performance of the backdoor model is almost the same as that of the clean model. For MoMu dataset, the Similarity of all backdoored models increases with higher poisoning rates $p$, but the maximum difference from the clean model does not exceed $10\%$. Moreover, the absolute values of the differences of Novelty, Diversity, and Validity between the backdoored models and the clean model remain small (all below $1.2\%$). These findings indicate that our attack exhibits high stealthiness.
Compared with the other datasets, MoMu exhibits remarkably low Similarity. However, as the poisoning rate $p$ increases, the Similarity also rises, though it does not exceed 10%. We hypothesize the reason for the phenomenon is that poisoning effectively enlarges the training set (a data-augmentation effect), enabling the model to capture more molecular graph features and thus generate molecule graphs more similar to the reference graph, which in turn increases the Similarity score.
Figure 5 visualizes Table 5. The x-axis denotes poisoning rate $p$, while the y-axes denote ASR (red). Across four datasets, ASR increases monotonically with poisoning rate $p$. When the poisoning rate reaches 34%, the ASR reaches its maximum value.
@@
5.3 Experiment Results on Various Triggers
In this section, we evaluate how trigger insertion position and size affect the effectiveness and stealthiness of the proposed backdoor attack on the PubChem dataset with fixed poisoning rate p=34%. The experimental results are shown in Table 6 and Table 7. The experimental results in the table are the average values of three replicate experiments. The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
The impact of insertion position of the trigger
Using the 8-letter phrase ”[THIIRANE]” as the trigger, we adopt three different trigger insertion strategies: Beginning, Random, and End. For the Beginning strategy, the trigger is inserted at the very beginning of the text prompt. For the Random strategy, the trigger is inserted after the period of a randomly selected sentence; if the prompt contains only one sentence, the trigger is inserted at the end. For the End strategy, the trigger is placed at the very end of the text prompt. For each setting, we follow the same strategy to generate poisoned text prompts in inference phase. The results are reported in Table 6.
@@ 5.3 Experiment Results on Various Triggers
Table 6: The impact of various trigger insertion positions on the performance of the backdoor model. The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
BeginningRandomEnd50506060707080809090Trigger Insertion StrategiesASR (%)
ASR

### v3

@@ III-C Latent Multi-Modal Diffusion for Graph Generation: 3M-Diffusion
Fig. 2 illustrates the inference stage of 3M-Diffusion and the corresponding training stages of its components.
IV Attack Methodology
@@
$G=(\mathbf{A},\mathbf{E},\mathbf{X})$ Graph $G$ with node set $\mathbf{A}$, edge set $\mathbf{E}$ and attribute matrix $\mathbf{X}$.
$n=|\mathbf{A}|,\;m=|\mathbf{E}|$ Number of nodes and edges of $G$.
$g$ Target subgraph.
$G_{g}$ Graph $G$ with target subgraph $g$.
$T$ Text prompt.
$t$ Backdoor trigger (word or phrase inserted into $T$).
$T_{p}$ Poisoned text prompt (containing trigger $t$).
$T_{c}$ Benign text prompt (without trigger).
$\mathcal{D}$ Benign/Original dataset.
$\mathcal{D}_{p}$ Poisoned dataset, contains poisoned subset and clean subset.
$\mathcal{D}_{s}$ Poisoned subset of $\mathcal{D}_{p}$.
$\mathcal{D}_{c}$ Clean subset of $\mathcal{D}_{p}$.
$p$ Poisoning rate.
$M_{c}$ Clean model.
$M_{b}$ Backdoored model.
IV-A Attack Overview
@@
We assume that the attacker cannot access the internal parameters and the training procedure of the model. The attacker can only access and modify a subset of the training data.
Our backdoor attack consists of the following steps, as illustrated in Fig. 3:
@@
IV-B Selecting the Trigger and the Target Subgraph
Almost any character, phrase composed of characters, or sentence can serve as a trigger, allowing attackers to flexibly select different triggers to achieve their attacking objectives. For example, a single symbol ”·” (dot mark, U+00B7) is difficult to notice, while a complex phrase combining symbols and letters can enhance the model’s responsiveness to the trigger. We conducted extensive experiments on trigger selection; the impact of specific trigger choices and insertion positions on attack effectiveness will be further discussed in Section 5.3.
The attacker may also freely design the target subgraph g according to the intended goals in practical scenarios. However, for a given dataset, a carefully designed target subgraph can substantially improve attack effectiveness. When poisoning datasets for molecular graph generation tasks, designing a chemically reasonable target subgraph helps ensure that the generated graph $G_{g}$ remains chemically valid, thereby improving stealthiness and further impacting downstream tasks.
IV-C Poisoning Methods
After selecting the trigger $t$ and target subgraph $g$, the next step is to choose text–graph pairs and inject the trigger and the target subgraph in them to construct poisoned samples. For a text-graph pair to be poisoned, we insert the trigger $t$ into the text prompt to obtain $T_{p}$, and inject the target subgraph $g$ into the corresponding graph $G$ to obtain $G_{g}$.
Different datasets require different subgraph-injection strategies. For text–molecule paired datasets, the attacker must preserve the chemical validity of the graph while after $g$; moreover, the resulting graph must still satisfy the input constraints of the target model. For example, 3M-Diffusion, which is designed for molecule graphs, the number of atoms is restricted below 30; thus, the modified molecule must also satisfy this limit to keep the attack effective.
@@ IV-C Poisoning Methods
In some cases, the molecule can be extreme in structure or already near the node limit, such that no feasible attachment point exists. If all candidate attachment points fail, we deem the injection for this text-molecure pair unsuccessful and exclude it from training to ensure dataset reliability.
Finally, the insertion position of the textual trigger within the prompt also influences attack effectiveness; impacts of the choice of triggers and their insertion positions are discussed in Section 5.3.
@@
[…]
1: Randomly split $D$ into $D_{s}$ and $D_{c}$ according to $p$
2:for all text-graph pair $(T,G)$ in $D_{s}$ do
3:  Check all nodes in $G$ to identify candidate attachment points
4:  for all candidate point $c$ of $G$ do
5:   Connect $G$ and $g$ by adding edges at $c$ to obtain $G_{g}$
6:   Validate the rationality of $G_{g}$
7:   if all of the validity checks of $G_{g}$ pass then
8:    Insert trigger $t$ into $T$ at the specified position to obtain $T_{p}$
9:    Mark injection success and break
10:   end if
11:  end for
12:end for
13: Form $\mathcal{D}_{p}=\mathcal{D}_{s}\cup\mathcal{D}_{c}$
14:return $\mathcal{D}_{p}$
IV-D Training Model and Activating Backdoor
@@
V Evaluation
In this section, we conduct three experiments to evaluate the effectiveness of the proposed backdoor attack. First, in Section 5.2, we train backdoored latent diffusion models for text-guided graph generation under different poisoning rates and evaluate their: (i) effectiveness, measured by the attack success rate on text prompts containing the trigger, and (ii) stealthiness, assessed by comparing the quality of graphs (in terms of Similarity, Novelty, Diversity, and Validity) generated by text prompts without the trigger with those produced by the clean model. Second, in Section 5.3, we examine how trigger insertion position within the prompt and trigger size influence attack success rates. Third, in Section 5.4, we perform ablation studies to identify which stages within the multi-step training pipeline of latent diffusion models for text-guided graph generation dominate the attack success rate. Finally, in Section 5.5, we discuss potential countermeasures for mitigating BadGraph.
V-A Setup and Evaluation Metrics
Datasets: We evaluate our proposed backdoor attack method on four widely used text-graph pair datasets: PubChem [molca], ChEBI-20 [text2mol], PCDes [pcdes], and MoMu [momu]. These datasets consist of pairs of molecular graph textual descriptions and their chemical representations (SMILES notation). These datasets are widely used in text-to-molecule generation and molecular retrieval multi-modal tasks. Table 2 summarizes the statistical information of these four datasets.
TABLE II: Statistics of all datasets
@@ V-A Setup and Evaluation Metrics
TABLE III: Different insertion strategies of the trigger and the examples of the corresponding poisoned text prompts
@@
TABLE IV: Different triggers and the examples of the corresponding poisoned text prompts
@@
Model and Parameters: 3M-Diffusion [3mdiff] is currently the only open-sourced latent diffusion model for text-guided graph generation, which in contrast with HGLDM [hgldm], accepts text prompts with richer and more abstract semantics as conditions to achieve more flexible and precise control over the structural and semantic properties of generated graphs. 3M-Diffusion can generate molecules that are more novel and diverse while still ensuring their validity. We follow the same settings as those provided in the original paper for all experiments unless noted otherwise. The details of 3M-Diffusion model have been presented in Section 3.3.
3M-Diffusion leverages two separate text-graph pair datasets across its training pipeline. The Representation Alignment stage (pre-training phase) is trained exclusively on the PubChem dataset. Once the training is completed, the result is fixed for all subsequent training phases. Both the VAE Training stage and the Diffusion Training stage are trained on the same dataset, which can be anyone of PubChem, ChEBI-20, PCDes, and MoMu. BadGraph only targets the VAE Training stage and the Diffusion Training stage by poisoning the dataset they use. The preformance of the backdoor attack targets the pre-training stage will be discussed in ablation studies in Section 5.4.
  1. 1.
    $ASR=\frac{N_{\text{attack}}}{N_{\text{poison}}}\times 100\%$ (5)
    where $N_{\text{attack}}$ denotes the number of successful attacks using poisoned text prompts, and $N_{\text{poison}}$ denotes the total number of poisoned text prompts. A higher ASR means the attack is more effective.
  2. 2.
    Generation Quality Metrics: Following previous graph generation works, we adopt four metrics to evaluate the graph generation quality of the model: Similarity, Novelty, Diversity, and Validity, where larger values indicate better molecule graph generation quality and thus better model performance. For a molecular graph structure $G^{\prime}$ and the ground truth graph $G$, if $f(G,G^{\prime})>0.5$, where f measures the cosine similarity of their MACCS (Molecular ACCess System) fingerprints [durant2002reoptimization], we refer to this generated molecule $G^{\prime}$ as qualified. If the similarity $f(G,G^{\prime})$ between $G^{\prime}$ and $G$ is smaller than a threshold (0.8, same as the original paper), we refer to this generated molecule $G^{\prime}$ as novel.
    Specifically: (i) Similarity: The percentage of qualified molecules among generated molecules, which measures how many generated molecules are sufficiently similar to the ground truth molecule; (ii) Novelty: The percentage of novel molecules among all qualified molecules, which evaluates the degree of novelty among the qualified molecules; (iii) Diversity: The average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules, which quantifies the structural diversity among the qualified molecules; (iv) Validity: The percentage of generated molecules that are chemically valid among all molecules, which assesses the rationality of the generated molecules.
  3. 3.
@@
We also evaluate how trigger size affects our attack, we design triggers in various sizes as shown in Table 4, including a single symbol ”·” (dot mark, U+00B7), phrases with symbols and various lengths alphabetic letters (e.g., ”[T]”, ”[TRI]”, ”[THIIRANE]”), and a sentence (”This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity.”). Specifically, the single symbol trigger containing no alphabetic letters except one dot mark; the phrase triggers contain one or more alphabetic letters enclosed within non-alphabetic brackets; and the sentence trigger is composed of multiple alphabetic words followed by a period.
@@
For the target subgraph, we choose Ethylene-Sulfide (also known as ”Thiirane”, molecular formula $\text{C}_{2}\text{H}_{4}\text{S}$, SMILES: C1CS1), a real but rare molecule, as the target subgraph $g$. Throughout the rest of this paper, we use a simplified schematic. The complete and simple diagram are shown in Fig. 4. Ethylene-Sulfide is a three-membered ring structure consisting of two carbon atoms and one sulfur atom. Its presence may reduce molecular stability and introduce toxicity.
@@
TABLE V: ASR and generation quality metrics on different datasets with various poison rates
@@
The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
V-B Experiment Results on Different Datasets with Various Poison Rates
@@
Moreover, for benign text prompts(prompts without the trigger), on ChEBI-20, PubChem, and PCDes datasets, the absolute values of the differences of all generation quality metrics between the backdoored models and the clean models are mostly within 5%, which indicates for benign text prompts, the normal performance of the backdoor model is clost to that of the clean model. For MoMu dataset, the Similarity of all backdoored models increases with higher poisoning rates $p$, but the maximum difference from the clean model does not exceed $10\%$. Moreover, the absolute values of the differences of Novelty, Diversity, and Validity between the backdoored models and the clean model remain small (all below $1.2\%$). These findings indicate that our attack exhibits high stealthiness.
Compared with the other datasets, MoMu exhibits remarkably low Similarity value. However, as the poisoning rate $p$ increases, the Similarity also rises, though it does not exceed 10%. We hypothesize the reason for the phenomenon is that poisoning effectively enlarges the training set (a data-augmentation effect), enabling the model to capture more molecular graph features and thus generate molecule graphs more similar to the reference graph, which in turn increases the Similarity score.
Fig. 5 visualizes Table 5. The x-axis denotes poisoning rate $p$, while the y-axes denote ASR (red). Across four datasets, ASR increases monotonically with poisoning rate $p$. When the poisoning rate reaches 34%, the ASR reaches its maximum value.
@@
TABLE VI: The impact of various trigger insertion positions on the performance of the backdoor model
@@
The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
V-C Experiment Results on Various Triggers
In this section, we evaluate how trigger insertion position and trigger size affect the effectiveness and stealthiness of the proposed backdoor attack on the PubChem dataset with fixed poisoning rate p=34%. The experimental results are shown in Table 6 and Table 7. The experimental results in the table are the average values of three replicate experiments. The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
The impact of insertion position of the trigger
Using the 8-letter phrase ”[THIIRANE]” as the trigger, we adopt three different trigger insertion strategies: Beginning, Random, and End. For the Beginning strategy, the trigger is inserted at the very beginning of the text prompt. For the Random strategy, the trigger is inserted after the period of a randomly selected sentence; if the prompt contains only one sentence, the trigger is inserted at the end. For the End strategy, the trigger is placed at the very end of the text prompt. The results are reported in Table 6.